OpenStudy (maddy1251):

Help? Will fan and medal? Can someone help me with these five problems, or at least the first one to help explain how to do these? For questions 1-5 label the equation as (a) an identity or (b) not an identity. 1. sin^2 x + cos^2 x = 1. (I think this is a not sure.) 2. sin x + cos x = 1 (is this not one? so b?) 3. sec x + csc x = 1 (I'm not sure) 4. cot^2 x = 1/tan^2 x (unsure..) 5. csc^2 x -1 = cot^2 x (Unsure..) If someone can help I'd really appreciate it.

OpenStudy (anonymous):

1) is definitely true, it is the mother of all trig identities

OpenStudy (anonymous):

2) is definitely not an identity you can check by plugging in some number for x and see that you don't get 1

OpenStudy (maddy1251):

@satellite73 would you mind explaining how that is proven? can you plug in a variable to get the answers?

OpenStudy (anonymous):

i think i can

OpenStudy (maddy1251):

Okay! Thank you SO MUCH.

OpenStudy (anonymous):

if you plug in a number and don't get the answer, then it is surely not an identity an identity means it is true for all x

OpenStudy (anonymous):

proving that it is an identity is a totally different story though

OpenStudy (maddy1251):

As in?

OpenStudy (anonymous):

are you used to working with degree, or radians?

OpenStudy (maddy1251):

A little

OpenStudy (anonymous):

i meant which ones do you use usually? \[\sin(\frac{\pi}{4}\]or \[\sin(45^\circ)\]

OpenStudy (maddy1251):

She has had us work with both, actually. I am just not a genius in it as I am in my advance algebra. Haha

OpenStudy (anonymous):

ok well then lets just say if you plug in a number (don't use 0 or 180) and you don't get the answer, then it is not an identity

OpenStudy (anonymous):

now how about \[\cot^2(x)=\frac{1}{\tan^2(x)}\]

OpenStudy (anonymous):

do you know the definition of cotangent?

OpenStudy (maddy1251):

not precisely.. all I know is it is A/O.

OpenStudy (anonymous):

ok that will work

OpenStudy (anonymous):

and tangent, to use your language, would be O/A

OpenStudy (maddy1251):

I remember that from SOA CAH TOA AH.

OpenStudy (anonymous):

i.e. cotangent is the reciprocal of tangent, and vice versa

OpenStudy (maddy1251):

That's the only way I have survived trig up to this point. I really don't like trig. And okay, gotcha!

OpenStudy (skullpatrol):

A simple algebraic example of an identity would be 4x + 10x = 14x. No matter what number you chose for x the LHS = RHS

OpenStudy (anonymous):

so \[\cot^2(x)=\frac{1}{\tan^2(x)}\] is just like saying \[\left(\frac{1}{z}\right)^2=\frac{1}{z^2}\] which is ture

OpenStudy (anonymous):

*true

OpenStudy (maddy1251):

I assume "z" is the varriable you used for cot and tan

OpenStudy (anonymous):

yes i was using z for tangent

OpenStudy (maddy1251):

Okay. It is basically showing that since tan is recipcrical to cot, that ratio you shows represents the fucntion between cot and tan, so that statement is then true. if that made sense.

OpenStudy (anonymous):

to write it all out, \[\cos^2(x)=\left(\frac{1}{\tan(x)}\right)^2=\frac{1}{\tan^2(x)}\]

OpenStudy (anonymous):

yes, what you said

OpenStudy (maddy1251):

Okay, I got you. Now what if there was no squared variables, would it still apply since cot is always equal to the recip of tan?

OpenStudy (anonymous):

i made a typo there, i meant cotangent, not cosine

OpenStudy (anonymous):

yes, tangent is the reciprocal of cotangent

OpenStudy (anonymous):

let me tell you what you should know besides " SOA CAH TOA"

OpenStudy (anonymous):

you should know that \[\tan(x)=\frac{\sin(x)}{\cos(x)}\] and \[\csc(x)=\frac{1}{\sin(x)}\] and \[\sec(x)=\frac{1}{\cos(x)}\]

OpenStudy (skullpatrol):

Without having to think

OpenStudy (anonymous):

using those identities, and \[\cos^2(x)+\sin^2(x)=1\] and algebra, you can solve most trig identities

OpenStudy (anonymous):

yes, without thinking just know them

OpenStudy (anonymous):

ready to tackle the last one?

OpenStudy (maddy1251):

I need to put them on index cards.

OpenStudy (skullpatrol):

Review them daily^

OpenStudy (maddy1251):

Yes I am. If I reply slow, it's ebcause OS is running slow for me now.

OpenStudy (anonymous):

i just sent you a decent cheat sheet, but yes, you should write them down yourself so you will know them

OpenStudy (anonymous):

yeah slow for me too, no problem

OpenStudy (maddy1251):

Thank you for the sheet!

OpenStudy (anonymous):

yw let me know when you are ready for the last one

OpenStudy (maddy1251):

\[\csc^{2}x-1=\cot^{2}x\]

OpenStudy (anonymous):

ok math teacher love this stuff this is where you start, with the mother of all identities,the first one \[\cos^2(x)+\sin^2(x)=1\]

OpenStudy (anonymous):

you don't have that, but you do have cosecant and cotangent do you remember what cosecant is? i wrote it above

OpenStudy (maddy1251):

you mean H/O or recip of sin?

OpenStudy (anonymous):

second one the reciprocal of sine

OpenStudy (maddy1251):

Okay.

OpenStudy (anonymous):

so starting with \[\cos^2(x)+\sin^2(x)=1\] how to you get \(\csc(x)\)? answer: divide everything by \(\sin(x)\) that way \(\csc(x)\) will turn up on the right

OpenStudy (maddy1251):

so it would be cos(x)/sin(x)

OpenStudy (maddy1251):

Or am I wrong

OpenStudy (anonymous):

\[\cos^2(x)+\sin^2(x)=1\\ \frac{\cos^2(x)}{\sin^2(x)}+\frac{\sin^2(x)}{\sin^2(x)}=\frac{1}{\sin^2(x)}\]

OpenStudy (anonymous):

i made a mistake, i should have said divide all by \(\sin^2(x)\)

OpenStudy (anonymous):

now you have almost exactly what you want

OpenStudy (anonymous):

\[\frac{1}{\sin^2(x)}=\csc^2(x)\] on the right

OpenStudy (anonymous):

evidently \[\frac{\sin^2(x)}{\sin^2(x)}=1\]

OpenStudy (maddy1251):

that comes from csc=1/sin correct? for the above? jsut to make sure

OpenStudy (anonymous):

yes what about \[\frac{\cos^2(x)}{\sin^2(x)}\] what is that ?

OpenStudy (maddy1251):

well.. cos/sin = cot, right? so cot^2(x)?

OpenStudy (anonymous):

zactly

OpenStudy (maddy1251):

Wow. I actually got that..

OpenStudy (anonymous):

so you have \[\cot^2(x)+1=\csc^2(x)\]

OpenStudy (anonymous):

yes, you did !

OpenStudy (maddy1251):

With your expert guidance of course.

OpenStudy (skullpatrol):

Can you think of any values that x can not have?

OpenStudy (anonymous):

you get from \[\cot^2(x)+1=\csc^2(x)\]to \[\csc^{2}x-1=\cot^{2}x\] using elementary algebra

OpenStudy (maddy1251):

Okay.

OpenStudy (anonymous):

aka add 1 to both sides

OpenStudy (anonymous):

ok not really, subtract 1 from both sides

OpenStudy (maddy1251):

I was gonna say you subtracted.

OpenStudy (maddy1251):

Dont lose me now! xD

OpenStudy (anonymous):

lol we are done in any case

OpenStudy (anonymous):

now i have a question: what the monkey is a "turkey committee"?

OpenStudy (maddy1251):

Haha, I think you have asked me before, I am almost positive. xD I am in 4-H and we show poultry, turkeys being one of the birds. There is committees that help run and function each species, including turkeys. I am part of the committee that helps assist other kids showing the project, taking care, being the advisor, etc.

OpenStudy (maddy1251):

I basically am like, co-boss.

OpenStudy (anonymous):

actually you are right , i remember now

OpenStudy (anonymous):

ok good luck with the trig, i am off to watch perry mason

OpenStudy (maddy1251):

Haha, okay, thank you so much for your help. You are my hero tonight.

OpenStudy (anonymous):

yw, my pleasure to help someone who is willing to help themselves

OpenStudy (maddy1251):

:) Take care!